OpenStudy (anonymous):

A conical tent made of canvas has a base that is 20 feet across and a slant height of 14 feet. To the nearest whole unit, what is the area of the canvas, including the floor? Use 3.14 for p. A. 754 ft2 B. 1,193 ft2 C. 2,135 ft2 D. 534 ft2

OpenStudy (nurali):

Area of Cone = pi*r*L + pi * r^2 where L is the slant height of the cone Surface Area = pi*10*14 + pi* (10)^2 Surface Area = 140 pi + 100 pi Surface Area = 240 pi Surface Area = 240*3.14 = 753.6 cubic feet approx